Solve the following boundary value problem.

uₙ = uₓₓ , 0 < x < L , n > 0

u(0,n) = u(L,n) = 0

u(x,0) = x(L-x) , 0 ≤ x ≤ L
1
Expert's answer
2019-08-28T05:14:08-0400

Let's assume that the solution will take the form

"u=\varphi(x)G(t)"

Substituting this from in the equation, will lead us to

"\varphi(x)G'(t)=\varphi''(x)G(t)"

"\frac{G'(t)}{G(t)}=\frac{\varphi''(x)}{\varphi(x)}"

Since the left hand side depends only of t and the right hand side only of x, the both sides can be equal only if they are equal to some constant.

Let's denote this constant as "-\lambda"

Then

"\frac{G'(t)}{G(t)}=\frac{\varphi''(x)}{\varphi(x)}=-\lambda"

Therefore we obtain two ordinary differential equations

"G'+\lambda G=0"

"\varphi''+\lambda\varphi=0"

Using boundary condition will lead us to

"\varphi(0)=\varphi(L)=0"

The general solution of the equation is

"\varphi=c_1\cos(\sqrt{\lambda}x)+c_2\sin(\sqrt{\lambda}x)"

Applying the first boundary condition gives us

"\varphi(0)=c_1=0"

Applying the second boundary condition, and using the above result, gives us

"\varphi(L)=c_2\sin(\sqrt{\lambda}L)=0"

We will only have a non-trivial solution, if "\sin(\sqrt{\lambda}L)=0" , "\lambda \not =0."

Hence

"\lambda_n=\left(\frac{\pi n}{L}\right)^2,\quad n\in \Zeta, n \not =0."

Then for G(t) we will have

"G'=-\lambda_nG"

This is a linear first order differential equation and its solution is

"G=B_ne^{-\lambda_nt }"

Hence the solution is

"u=\sum\limits_{n=1}^{\infty}B_ne^{-\lambda_nt }\sin\sqrt{\lambda}x"

The initial condition will lead us to

"u(x,0)=\sum\limits_{n=1}^{\infty}B_n\sin\sqrt{\lambda}x=x(L-x)"

which is the sine Fourier series of "x(L-x)"


It's coefficients can be obtained using the formula

"B_n= \frac{2}{L}\int\limits_{0}^{L}x(L-x)sin(\frac{\pi n}{L})xdx=-\frac{2 L^2 (-2 + 2 cos(n π) + n π sin(n π))}{n^3 π^3}=-\frac{4 L^2((-1)^n-1)}{n^3 π^3}"

Hence

"B_{2n}=0"

"B_{2n-1}=\frac{8 L^2}{(2n-1)^3 π^3}"

Therefore we obtain the final form of the function

"u=\sum\limits_{n=1}^{\infty}\frac{8 L^2}{(2n-1)^3 π^3}e^{-\lambda_{2n-1}t }\sin\sqrt{\lambda_{2n-1}}x"